# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ D.2. Baseline Methods
  • SAGE: This algorithm adopts the graph perturbation, building on the popular GraphSAGE architecture as its backbone GNN model. To realize graph perturbation, SAGE perturbs the adjacency matrix of graph using the asymmetric randomized response.
@@
Vanilla algorithms.
@@
The non-private versions of all private algorithms above are used to quantify the accuracy loss of corresponding EDP algorithms.
Appendix E More Experimental Results
E.1. Computational Overhead
Table 7. Execution Time of $3$ Training-Test Runs (seconds)
@@ E.1. Computational Overhead
Dataset $\mathsf{CARIBOU}$ DPDGC GAP SAGE MLP Chain-S $2.66$ $3.70$ $5.15$ $2.97$ $2.34$ Chain-M $2.66$ $3.70$ $5.24$ $2.98$ $2.34$ Chain-L $2.65$ $3.69$ $5.31$ $3.00$ $2.35$ Chain-X $2.67$ $3.73$ $5.27$ $2.97$ $2.36$ Computers $3.34$ $4.47$ $6.19$ $3.41$ $2.56$ Facebook $4.01$ $6.04$ $6.30$ $3.74$ $2.74$ PubMed $3.10$ $4.63$ $6.32$ $3.34$ $2.52$ Cora $2.79$ $4.02$ $5.52$ $3.04$ $2.40$ Photo $3.04$ $3.97$ $5.93$ $3.18$ $2.48$
We measure the latency and memory usage to assess the overhead of $\mathsf{CARIBOU}$. For latency, we use the total execution time (training and testing), and in Table 7 we list the average of running $3$ times. We found $\mathsf{CARIBOU}$ has less latency than the other method (2.99 seconds of $\mathsf{CARIBOU}$ averaged among datasets, comparing to 4.22 seconds of DPDGC, 5.69 seconds of GAP, 3.18 seconds of SAGE), except MLP (2.45 seconds), which does not incur any graph operations. Notable latency increase is observed on the Facebook dataset because its large number of nodes (26,406) and edges (2,117,924), and $\mathsf{CARIBOU}$ shows a larger lead comparing the other private GNNs: 4.01, 6.04 and 6.30 for $\mathsf{CARIBOU}$, GAP and p̣ d g c.
For memory, we measure the max memory usage over the 3 runs and report the results in Table 8. For chain-structured datasets, the memory consumption averaged across datasets of $\mathsf{CARIBOU}$ (1.09MB) is close to SAGE (1.12MB) and MLP (1.05MB), while much lower than GAP (1.55MB) and DPDGC (1.55MB). On the other hand, we found higher memory consumption of $\mathsf{CARIBOU}$ is introduced on the other datasets, comparing to the baseline methods due to graph loading mechanism in GAP. Still, the memory consumption of $\mathsf{CARIBOU}$ is reasonable, which is smaller than our GPU memory limit (24GB) for all datasets. We argue the privacy, and utility benefits of $\mathsf{CARIBOU}$ outweigh its memory costs here.
Table 8. Max Memory Usage (MB)
@@
Dataset $\mathsf{CARIBOU}$ DPDGC GAP SAGE MLP Chain-S $1.08$ $1.46$ $1.39$ $1.09$ $1.05$ Chain-M $1.08$ $1.49$ $1.44$ $1.10$ $1.05$ Chain-L $1.09$ $1.56$ $1.56$ $1.12$ $1.06$ Chain-X $1.10$ $1.68$ $1.80$ $1.16$ $1.07$ Computers $314.04$ $197.21$ $138.39$ $116.85$ $134.13$ Facebook $525.18$ $494.07$ $315.82$ $268.41$ $275.01$ PubMed $272.70$ $202.67$ $190.10$ $97.45$ $111.17$ Cora $104.94$ $56.14$ $27.53$ $32.99$ $42.19$ Photo $170.28$ $110.60$ $77.54$ $62.69$ $72.29$
E.2. Accuracy of $\mathsf{CARIBOU}$-NDP
@@
Table 10 presents the top accuracy for NDP implemented by $\mathsf{CARIBOU}$ on 6 datasets. The results demonstrate that accuracy increases in general (except Chain-S at $\epsilon=16,32$) with larger $\epsilon$ values, aligning with the EDP results shown in Table 3. As established in Theorem 9, NDP requires injecting more noise compared to EDP under the same privacy budget, hence, the accuracy of NDP is often lower than EDP.
Regarding the other baselines, we we did not implement the NDP version of GAP because it does not correctly account for node-level sensitivity, as pointed out by Xiang et al. (Xiang et al., 2024). Other baselines like DPDGC do not aim at multi-hop GNN training, i.e., with a fixed number of layers, thus, we do not test the NDP versions for them.
@@
Table 9. Mean accuracy over 3 runs for EDP. The best accuracy and the second-best accuracy are highlighted, respectively. The symbol

### v2

@@ A-D3 Baseline algorithms
  • PertGraph: This algorithm adopts the graph perturbation, building on the popular GraphSAGE architecture as its backbone GNN model. To realize graph perturbation, PertGraph perturbs the adjacency matrix of graph using the asymmetric randomized response.
@@
By extending edge-level private algorithms, node-level DP detailed below can be realized by protecting full information (i.e., edges, node features) of a node.
  • DPDGC: This algorithm extends DPDGC and bounds the out-degree of nodes to realize the node-level DP. Thus, DPDGC reduces the dependency of DP noise variance on the maximum node degree, improving the trade-off between privacy guarantee and model utility.
  • GAP: This algorithm extends GAP for bounded-degree graphs, where each node has controllable influence to its neighbors by sampling a limited number of neighbors.
  • PertGraph: This algorithm adapts DP-SGD to the GraphSAGE model, and simultaneously adds the noise to aggregation function constrained by node-level sensitivity.
  • MLP: This algorithm is trained with DP-SGD without accessing the edges.
A-D4 Vanilla algorithms
@@
The non-private versions of all private algorithms above are used to quantify the accuracy loss of corresponding EDP and NDP algorithms.
A-D5 Model architectures & configurations of $\mathsf{CARIBOU}$
We concatenate $K$ layers ranging from $1$ to $20$, with $K,\alpha_{1},\beta,C_{\mathsf{L}}$ as the hyper-parameters. If the range between two interactive nodes is long, we use larger values to test the influence of hops. Specifically, for Chain-S and Chain-M datasets, $\mathsf{CARIBOU}$ uses the number of hops $K\in\{15,10,9,8,7,6,5,4,3,2,1\}$, while for Chain-S and Chain-M datasets consisting of longer chains, $\mathsf{CARIBOU}$ takes $K\in\{20,17,15,\allowbreak 13,11,9,7,5\}$. As for standard datasets listed in Table V, $K$ is chosen from the union sets of the aforementioned two sets, i.e., $\{20,17,15,\allowbreak 13,11,10,9,8,7,6,5,4,3,2,1\}$. To align with GAP’s configuration [uss/sajadmanesh2023gap], we set the number of hidden units to $\{16,64\}$ and use the SeLU activation function [nips/KlambauerUMH17] at every layer. We adopt the Adam optimizer over $100$ epochs with a learning rate $0.001$, and pick the best accuracy to report.
A-D6 Privacy configuration and parameters
We implement convergent privacy allocation based on GAP’s privacy budget accounting mechanism [uss/sajadmanesh2023gap] and numerically calibrate noise level $\sigma$ by setting $\epsilon$. All edge/node-level private algorithms adopt the Gaussian mechanism to sample the i.i.d. noise to realize the desired $(\epsilon,\delta)$-DP. For comprehensive evaluation, we consider different choices of $\epsilon\in\{1,2,4,8,16,32\}$, while $\delta$ is set to be smaller than the inverse number of edges for EDP or the inverse number of nodes for NDP.
A-E Computational Overhead
TABLE VII: Execution Time of 33 Training-Test Runs (seconds) for EDP
@@ A-E Computational Overhead
Dataset $\mathsf{CARIBOU}$ DPDGC GAP PertGraph MLP Computers $3.70$ $87.72$ $6.62$ $3.37$ $39.37$ Facebook $4.34$ $233.65$ $6.36$ $4.29$ $70.93$ PubMed $3.46$ $149.07$ $6.08$ $3.78$ $55.11$ Cora $3.05$ $16.85$ $5.98$ $3.10$ $10.88$ Photo $3.07$ $43.84$ $6.00$ $3.33$ $23.96$ Chain-S $2.92$ $5.93$ $5.54$ $2.99$ $4.85$ Chain-M $2.76$ $5.62$ $5.67$ $3.16$ $4.76$ Chain-L $2.77$ $5.93$ $5.39$ $3.15$ $4.61$ Chain-X $2.91$ $5.93$ $5.23$ $3.06$ $5.04$
TABLE VIII: Execution Time of 33 Training-Test Runs (seconds) for NDP
@@
Dataset $\mathsf{CARIBOU}$ DPDGC GAP PertGraph MLP Computers $3.23$ $526.12$ $351.83$ $386.04$ $1452.91$ Facebook $4.39$ $2056.61$ $726.72$ $767.28$ $2217.46$ PubMed $2.82$ $1183.27$ $528.56$ $418.21$ $1472.64$ Cora $2.52$ $96.29$ $92.04$ $62.06$ $184.39$ Photo $2.91$ $298.93$ $217.79$ $203.94$ $329.01$ Chain-S $2.26$ $10.96$ $9.41$ $1.51$ $13.77$ Chain-M $2.27$ $11.37$ $10.52$ $1.49$ $13.56$ Chain-L $2.29$ $10.48$ $7.71$ $1.53$ $13.74$ Chain-X $2.26$ $11.66$ $7.21$ $1.51$ $13.63$
We measure the latency and memory usage to assess the overhead of $\mathsf{CARIBOU}$ when $\epsilon=1,K=20$ as an example.
A-E1 Execution time
For latency, we use the total execution time (training and testing) for EDP, and in Table VII we list the average of running $3$ times. We found $\mathsf{CARIBOU}$ has less latency than baselines including $3.22$ seconds of $\mathsf{CARIBOU}$ averaged among datasets, comparing to $61.8$ seconds of DPDGC, $5.99$ seconds of GAP, and $24.39$ seconds MLP, except for comparative $3.25$ seconds of PertGraph. Notable latency increase is observed on the Facebook dataset because its large number of nodes (26,406) and edges (2,117,924), and $\mathsf{CARIBOU}$ shows a larger lead comparing the other private GNNs: $233.65$ and $6.36$ for DPDGC and GAP, respectively.
As for NDP, among all the evaluated models, $\mathsf{CARIBOU}$ consistently achieves the lowest execution times across all datasets, with an average of only $2.77$ seconds per training-test run. This is substantially faster than the other methods, particularly in comparison to resource-intensive models like DPDGC and MLP, which have average execution times of $467.52$ seconds and $633.57$ seconds, respectively. $\mathsf{CARIBOU}$’s ability to process both large and small datasets demonstrates its potential advantage for real-world applications where computational resources or time may be limited.
A-E2 Memory costs
For memory, we measure the max memory usage over the 3 runs and report the results in Table IX under EDP. For chain-structured datasets, the memory consumption averaged across datasets of $\mathsf{CARIBOU}$ (1.09MB) is close to PertGraph (1.12MB) and MLP (1.05MB), while much lower than GAP (1.55MB) and DPDGC (1.55MB). On the other hand, we found that a higher memory consumption of $\mathsf{CARIBOU}$ is introduced in the standard graph datasets, compared to the baseline methods due to the graph loading mechanism in GAPand controllable parameters for contractiveness. Still, the memory consumption of $\mathsf{CARIBOU}$ is reasonable, which is smaller than our GPU memory limit (24GB) for all datasets. We argue the privacy, and utility benefits of $\mathsf{CARIBOU}$ outweigh its memory costs here.
TABLE IX: Max Memory Usage (MB) for EDP
Dataset $\mathsf{CARIBOU}$ DPDGC GAP PertGraph MLP Computers $314.04$ $1833.63$ $138.40$ $154.57$ $129.27$ Facebook $525.18$ $3657.00$ $315.84$ $359.63$ $251.24$ PubMed $272.70$ $2594.26$ $190.12$ $167.64$ $98.26$ Cora $104.94$ $377.78$ $27.55$ $32.85$ $77.01$ Photo $170.28$ $1021.70$ $77.54$ $83.49$ $80.14$ Chain-S $1.08$ $1.34$ $1.41$ $1.10$ $1.03$ Chain-M $1.09$ $1.42$ $1.46$ $1.11$ $1.03$ Chain-L $1.09$ $1.69$ $1.58$ $1.13$ $1.03$ Chain-X $1.11$ $2.42$ $1.82$ $1.17$ $1.04$
TABLE X: Max Memory Usage (MB) for NDP. Max node degree is 20.
Dataset $\mathsf{CARIBOU}$ DPDGC GAP PertGraph MLP Computers $357.28$ $1831.35$ $319.80$ $1621.09$ $130.88$ Facebook $531.76$ $3484.93$ $479.18$ $1676.24$ $223.17$ PubMed $314.45$ $2597.45$ $345.10$ $577.42$ $101.49$ Cora $121.39$ $469.11$ $299.51$ $890.55$ $78.15$ Photo $195.02$ $1051.85$ $239.85$ $1352.86$ $81.29$ Chain-S $1.40$ $1.37$ $2.13$ $1.10$ $1.09$ Chain-M $1.41$ $1.44$ $2.32$ $1.11$ $1.10$ Chain-L $1.42$ $1.71$ $2.79$ $1.13$ $1.11$ Chain-X $1.46$ $2.44$ $3.72$ $1.19$ $1.13$
TABLE XI: Privacy Auditing via LinkTeller and G-MIA. AUC score is reported.
Dataset LinkTeller G-MIA $\epsilon=\inf$ $\epsilon=1$ $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=16$ $\epsilon=32$ $\epsilon=\inf$ $\epsilon=1$ $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=16$ $\epsilon=32$ Facebook $0.977$ $0.463$ $0.483$ $0.482$ $0.478$ $0.472$ $0.462$ $0.567$ $0.587$ $0.587$ $0.587$ $0.583$ $0.587$ $0.588$ PubMed $0.981$ $0.446$ $0.442$ $0.441$ $0.445$ $0.442$ $0.449$ $0.600$ $0.599$ $0.598$ $0.598$ $0.601$ $0.601$ $0.605$ Cora $0.998$ $0.427$ $0.448$ $0.399$ $0.443$ $0.451$ $0.450$ $0.645$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ Photo $0.962$ $0.475$ $0.404$ $0.421$ $0.428$ $0.417$ $0.434$ $0.678$ $0.677$ $0.682$ $0.682$ $0.678$ $0.672$ $0.676$ Computers $0.860$ $0.367$ $0.364$ $0.372$ $0.363$ $0.361$ $0.384$ $0.702$ $0.701$ $0.661$ $0.707$ $0.708$ $0.711$ $0.701$
Table X presents the maximum memory usage (in MB) for NDP across a variety of datasets with a maximum node degree of 20, highlighting the efficiency of our method, $\mathsf{CARIBOU}$. Across all datasets, $\mathsf{CARIBOU}$ demonstrates significantly lower memory consumption compared to DPDGC, which incurs the highest memory usage, particularly on large-scale datasets like Facebook (3484.93 MB for DPDGC vs. 531.75 MB for $\mathsf{CARIBOU}$) and PubMed (2597.45 MB for DPDGC vs. 314.44 MB for $\mathsf{CARIBOU}$). While baseline methods such as GAP and MLP show lower memory usage on certain datasets, $\mathsf{CARIBOU}$ achieves a favorable balance of efficiency and scalability, maintaining competitive memory consumption across both small and large graphs. Notably, on the Cora and Photo datasets, $\mathsf{CARIBOU}$ uses 104.51 MB and 165.86 MB, respectively, which is substantially less than DPDGCand only moderately higher than the most memory-efficient baselines. Notably, PertGraph incurs much larger memory overhead under NDP than that of EDP settings due to its $O(|{\bm{V}}|^{2})$ computational complexity. Overall, these results illustrate that $\mathsf{CARIBOU}$ effectively controls memory usage, outperforming DPDGC by a wide margin and providing robust scalability under NDP settings.
A-F More Results of EDP and NDP
A-F1 Analysis of EDP
Regarding the standard graph datasets, for Computers, PubMed, Cora and Photo, $\mathsf{CARIBOU}$ can outperform all the other baselines in most cases with varying $\epsilon$. It is also worth noting that $\mathsf{CARIBOU}$ achieves fairly stable accuracy with with varying $\epsilon$ (e.g., $92.0\%$ to $92.4\%$ for Computers with $\epsilon$ growing from $1$ to $\inf$), suggesting the convergent privacy design is indeed effective. For Facebook, we found $\mathsf{CARIBOU}$ outperforms the other baselines with small $\epsilon$ (i.e., $\epsilon\leq 4$), but falls behind at larger $\epsilon$. Yet, we argue that performance at low privacy budget is more critical for privacy-sensitive datasets like Facebook, and stable accuracy across privacy budget $\epsilon$ might be more desired.
Now, we highlight the results on the chain-structured datasets. In summary, we observe that the non-private versions (i.e., $\epsilon=\inf$) of GAP and $\mathsf{CARIBOU}$ can achieve perfect classification accuracy, while DPDGC, PertGraph and MLP show relatively low accuracy. The root cause is that features of a distant node can only be learned through a sufficient number of hops of message passing, i.e., aggregating and passing features of the neighboring node one by one, but DPDGC and PertGraph are not designed to leverage long-hop interactions. For MLP, as the graph structure is not utilized at all, the result is close to random guessing. After adding DP noises to protect edge-level memberships, the accuracy of both GAP and $\mathsf{CARIBOU}$ drop, but the accuracy loss of GAP is more prominent (from $100\%$ down to $57.8\%$-$62.5\%$). GAP is able to maintain reasonable accuracy even with fairly tight privacy budget $\epsilon=1$ on small chain-structured datasets like Chain-S and Chain-M. On the other hand, Chain-L and Chain-X have relatively low accuracy even for $\mathsf{CARIBOU}$ (e.g., $70.0\%$ and $66.0\%$ when $\epsilon=1$). This observed phenomenon is attributed to the increased difficulty of GNN classification with longer chains, as DP noise is injected at every layer, compounding its impact.
A-F2 Analysis of NDP
@@
In line with the EDP results, the NDP results also demonstrate that precision increases in general with larger $\epsilon$ values. As established in Theorem 5, NDP requires injecting more noise compared to EDP under the same privacy budget, hence, the accuracy of NDP is often lower than EDP for standard datasets. This phenomenon does not hold for chain-structured datasets as the original max node degree is very small, i.e., $0,1,2$. Ablation study on configuring different max node degree for NDP is in Section VI-C5.
Notably, $\mathsf{CARIBOU}$ consistently outperforms all other methods across all privacy budgets $\epsilon$ and across all nine datasets, while DPDGC shows the second-best performance overall but remains significantly behind $\mathsf{CARIBOU}$. For instance, $\mathsf{CARIBOU}$ achieves $91.91\%$, which is $35\%$+ higher than the next best MLP of $63.44\%$, $54\%$ higher than DPDGC of $56.72\%$, and more than $60\%$ higher than PertGraphwith $29.51\%$. MLPrepresents the baseline of learning node features independently without graph topology. In particular, $\mathsf{CARIBOU}$ is the only framework that can surpass MLPin many cases, showing effective GNN learning over structural graphs. In addition, GAP, PertGraph, and MLP generally yield lower values as the realization of NDP is more challenging than the that of EDP. Thus, Table III indicates strong and stable advantage of $\mathsf{CARIBOU}$ over state-of-the-art baselines in NDP settings.
A-F3 More Results
Table XIII shows the mean accuracy compared between $\mathsf{CARIBOU}$ and other baselines across all datasets. For both EDP and NDP tasks, $\mathsf{CARIBOU}$ always achieves the highest or second-highest accuracy, often outperforming competing methods by a notable margin. This is especially apparent in the NDP task, where the performance gap widens under stricter privacy budgets (lower $\epsilon$), demonstrating $\mathsf{CARIBOU}$’s robustness in privacy-sensitive regimes. For example, under NDP with $\epsilon=1$ on the Cora dataset, $\mathsf{CARIBOU}$ achieves an accuracy of $67.47\%$, while the next best method, DPDGC, reaches just $31.04\%$ with a difference of over $50\%$ drop. In the chain-structured datasets, $\mathsf{CARIBOU}$ maintains strong performance even as the privacy constraints increase, whereas other models often see substantial accuracy drops. Notably, in standard graph datasets and higher privacy budgets, $\mathsf{CARIBOU}$’s accuracy advantage is frequently greater than $10\%$ compared to the next best approach, as indicated by green arrows in the table. Even in the non-private setting, $\mathsf{CARIBOU}$ delivers near-optimal or best accuracy across all datasets. These results highlight $\mathsf{CARIBOU}$’s overall reliability, adaptability, and superior utility for both private and non-private graph learning tasks.
Figure 7: Accuracy of non-private $\mathsf{CARIBOU}$ and GAP under different $K$.
Figure 8: Ablation study of max node accuracy for NDP.
Figure 9: Ablation Study of $K$ on $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines). For each pair of $\epsilon$ and dataset (Chain-S, Chain-M and Chain-L), different $K$ are used.
Figure 10: Ablation Study of $\epsilon$ on $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines). For each pair of $K$ and dataset (Chain-S, Chain-M and Chain-L), different $\epsilon$ are used.
@@ A-F3 More Results
TABLE XII: Top Accuracy over 3 Runs for NDP. Maximum node degree is 10 for relatively large datasets (PubMed, Facebook) and 5 for other datasets. The best accuracy and the second-best accuracy are highlighted, respectively.
Dataset Computers Facebook PubMed Cora Photo Chain-S Chain-M Chain-L Chain-X $\epsilon=1$ $\mathsf{CARIBOU}$ $91.91\%$ $60.36\%$ $73.85\%$ $80.63\%$ $95.10\%$ $78.12\%$ $70.0\%$ $61.67\%$ $61.0\%$ DPDGC $57.65\%$ $40.93\%$ $58.91\%$ $33.21\%$ $42.21\%$ $58.06\%$ $57.5\%$ $57.63\%$ $54.6\%$ GAP $36.71\%$ $35.66\%$ $53.06\%$ $34.13\%$ $32.07\%$ $65.62\%$ $55.0\%$ $58.33\%$ $59.0\%$ PertGraph $34.41\%$ $20.49\%$ $39.89\%$ $1.03\% 22$$2.8\%$ $59.38\%$ $60.0\%$ $6.67\% 55$$5.0\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $92.28\%$ $64.95\%$ $80.4\%$ $83.76\%$ $94.76\%$ $78.12\%$ $72.5\%$ $63.33\%$ $61.0\%$ DPDGC $66.11\%$ $46.95\%$ $70.94\%$ $31.37\%$ $53.41\%$ $58.06\%$ $57.5\%$ $57.63\%$ $54.5\%$ GAP $47.48\%$ $39.80\%$ $68.93\%$ $32.47\%$ $37.31\%$ $65.62\%$ $55.0\%$ $58.33\%$ $59.0\%$ PertGraph $35.63\%$ $20.33\%$ $40.02\%$ $24.35\%$ $23.92\%$ $59.38\%$ $60.0\%$ $6.67\% 55$$5.0\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $92.28\%$ $68.19\%$ $84.81\%$ $85.98\%$ $94.9\%$ $78.12\%$ $72.5\%$ $63.33\%$ $62.0\%$ DPDGC $72.20\%$ $48.77\%$ $80.02\%$ $32.66\%$ $69.91\%$ $58.06\%$ $57.5\%$ $57.63\%$ $59.6\%$ GAP $61.77\%$ $46.72\%$ $79.51\%$ $33.58\%$ $45.79\%$ $65.62\%$ $55.0\%$ $58.33\%$ $59.0\%$ PertGraph $35.56\%$ $21.09\%$ $40.78\%$ $26.57\%$ $25.38\%$ $59.38\%$ $60.0\%$ $6.67\% 55$$5.0\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $92.24\%$ $70.11\%$ $87.34\%$ $87.45\%$ $94.96\%$ $75.00\%$ $72.5\%$ $66.67\%$ $63.0\%$ DPDGC $76.39\%$ $49.91\%$ $83.49\%$ $43.91\%$ $77.67\%$ $58.06\%$ $57.5\%$ $57.63\%$ $58.6\%$ GAP $68.52\%$ $48.24\%$ $82.17\%$ $31.73\%$ $68.39\%$ $65.62\%$ $55.0\%$ $58.33\%$ $59.0\%$ PertGraph $34.41\%$ $22.25\%$ $42.28\%$ $28.97\%$ $27.17\%$ $59.38\%$ $60.0\%$ $6.67\% 55$$5.0\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $92.50\%$ $70.60\%$ $88.49\%$ $87.82\%$ $94.96\%$ $71.88\%$ $72.5\%$ $73.33\%$ $65.0\%$ DPDGC $78.88\%$ $50.11\%$ $84.88\%$ $56.46\%$ $83.10\%$ $58.06\%$ $57.5\%$ $57.63\%$ $54.5\%$ GAP $73.83\%$ $49.91\%$ $83.77\%$ $37.27\%$ $76.61\%$ $65.62\%$ $55.0\%$ $58.33\%$ $59.0\%$ PertGraph $33.18\%$ $22.15\%$ $44.00\%$ $30.44\%$ $29.62\%$ $59.38\%$ $60.0\%$ $6.67\% 55$$5.0\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $92.46\%$ $70.96\%$ $88.97\%$ $87.82\%$ $94.90\%$ $68.75\%$ $72.5\%$ $75.00\%$ $64.0\%$ DPDGC $81.40\%$ $50.64\%$ $86.33\%$ $64.21\%$ $86.08\%$ $58.06\%$ $57.5\%$ $57.63\%$ $54.5\%$ GAP $77.13\%$ $50.66\%$ $85.21\%$ $57.38\%$ $80.38\%$ $62.50\%$ $55.0\%$ $56.67\%$ $59.0\%$ PertGraph $36.45\%$ $22.67\%$ $46.97\%$ $31.73\%$ $34.13\%$ $59.38\%$ $60.0\%$ $6.67\% 55$$5.0\%$
TABLE XIII: Mean Accuracy over 3 Runs for EDP and NDP. The best accuracy and the second-best accuracy are highlighted, respectively. The symbol